OpenStudy (just_gymnast_me):

Please help! Will fan and medal!!! Solve the proportion x over x minus 4 equals 4 over x minus 4.. What is the value of x? x = 2 x = −2 and x = 1 x = −1 and x = 4 x = 4

OpenStudy (just_gymnast_me):

\[\frac{ x }{ x-4 } = \frac{ 4 }{ x-4 }\]

OpenStudy (just_gymnast_me):

im sure the first x in the numerator is 4 then cross multiply??

OpenStudy (anonymous):

Multiply both sides by x-4 and get x=4.